From the top of an 80-ft. building, the angle of elevation of the top of a taller building is 49 and

the angle of depression of the base of this building is 62. Determine the height of the taller

building to the nearest foot.


62°

49°


80 ft.


a. 211 ft. b. 112 ft. c. 129 ft. d. 276 ft.


1
Expert's answer
2021-06-03T12:18:44-0400

1) Firstly we should find the distance between two buildings. It can be done using the angle of depression, 62. We can imagine a triangle ABC formed by the distance between the buildings, AB = x, the height of the shorter building, BC = 80ft, which is projected to the taller building, and the distance between the bottom of the higher building and the top of the lower one, CA. The angle of depression ∠ CAB = 62, this triangle is right, so we can find the ∠ ACB to use it for the law of sines.

"\u2220 ACB = 180\u00b0 - \u2220 ABC - \u2220 CAB = 180\u00b0 - 90\u00b0 - 62\u00b0 = 28\u00b0"

The law of sines is an equation relating the lengths of the sides of a triangle to the sines of its angles.


"a\/sin A=b\/sin B = c\/Sin C=2R"

It helps to find a side of a triangle with 2 angles and other side which are known.



For the present task:


"sin(\u2220ACB)\/BA=sin(\u2220CAB)\/BC"

"AB=sin(\u2220ACB)*BC\/sin(\u2220CAB)= sin(28\u00b0)*80\/sin(62\u00b0)="


"=0.469*80\/0.883=42.5 (ft)"

Then image another triangle, ABD, where D is the top of the higher building. Using the same law of sines and the angle of elevation, 49, we can make the following proportion:

"sin(\u2220ADB)\/BA=sin(\u2220DAB)\/BD"


"BD=sin(\u2220DAB)*AB\/sin(\u2220ADB)"

"BD = sin(49)*42.5\/sin(180-90-49)=0.755*42.5\/0.656 = 48.72"

Adding two parts of the taller building, we have:


"CD = CB + BD = 80 + 48 = 128.72 (ft)"

Rounding to the nearest foot, we have "CD \u2248 129 ft."

Answer: the height of the taller building is 129 ft, c.